40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Aug 18 23:59:53 2017, gautam, Update, ALS, X Arm ALS lock 
    Reply  Sat Aug 19 01:35:08 2017, ericq, Update, ALS, X Arm ALS lock ALS_comparison.pdf
    Reply  Mon Aug 21 23:38:55 2017, gautam, Update, ALS, ALS out-of-loop noise ALS_21082017.pdf
       Reply  Tue Aug 22 02:19:11 2017, gautam, Update, ALS, ALS OLTFs ALS_OLTFs.pdfALS_OLTF_Aug2017.zip
          Reply  Tue Aug 22 23:27:14 2017, rana, Update, ALS, ALS OLTFs 
Message ID: 13229     Entry time: Fri Aug 18 23:59:53 2017     Reply to this: 13230   13237
Author: gautam 
Type: Update 
Category: ALS 
Subject: X Arm ALS lock 

[ericq, gautam]

  • I was just getting the IFO aligned, and single arm lock going, when EricQ came in and asked if we could get some ALS data.
  • ALS beats seemed fine, in particular the X-Arm. The broad hump around ~70Hz that was present in my previous ALS update was nowhere to be seen - reasons unknown.
  • Copied over /opt/rtcds/caltech/c1/scripts/YARM/Lock_ALS_YARM.py to /opt/rtcds/caltech/c1/scripts/XARM/Lock_ALS_XARM.py. Could be useful when we want to do arm cavity scans.
  • Made appropriate changes to allow ALS locking of Xarm - the testpoint inaccessibility makes things a little annoying but for tonight we just used DQ channels in place (or slow channels when DQ chans were not available)
  • Calibration of X arm error signal seemed off - so we fixed it by driving a line in ETMX and matching up the peaks in the ALS error signal and POX11. We then updated the gain of the filter in the CINV filter bank accordingly.
  • Got some decent data - X arm stayed locked on ALS for >60mins, during which time the Y arm stayed locked on POY11, and the Y green also reained locked yes. There was no evidence of the X arm 00 mode randomly dropping out of lock tonight.
  • EQ will update with a sick comparison plot - today we looked at the ALS noise from the perspective of the Green Locking Izumi et. al. paper.
  • Y arm ALS noise didn't look so hot tonight - to be investigated...

Leaving LSC mode OFF for now while CDS is still under investigation


Not really related to this work: We saw that the safe.snap file for c1oaf seems to have gotten overwritten at some point. I restored the EPICS values from a known good time, and over-wrote the safe.snap file.

ELOG V3.1.3-